# Engineering Software 2.0 by Interpolating Neural Networks: Unifying Training, Solving, and Calibration

### v3

@@ 2 INN: Unifying Training/Learning, Solving, and Calibration
A general INN forward propagation is provided in Eq.3 as a tensor contraction. The INN trainer optimizes nodal values $\bm{U}=\{\bm{u}^{(j)}\}_{j=1,\cdots,J}$ (and if needed, nodal coordinates ($\bm{X}=\{\bm{x}^{(j)}\}_{j=1,\cdots,J}$)) under a given loss function and training data. Consider a regression problem with $K$ labeled data: $(\bm{x}^{*}_{k},\bm{u}^{*}_{k})$, $k=1,\cdots,K$ and $\bm{x}^{*}_{k}\in\mathbb{R}^{I},\bm{u}^{*}_{k}\in\mathbb{R}^{L}$. Here the superscript * denotes data. A mean squared error (MSE) loss function for this regression problem is defined as:
@@
Finally, an optimization problem is formulated as:
@@
Several benchmark problems are introduced in the Supplementary Information. The superior training accuracy, faster training and inference time, and fewer trainable parameters of the INN trainer are demonstrated.
@@
INN solver generalizes classical numerical methods such as FEM and meshfree, as well as the model order reduction methods such as proper generalized decomposition (PGD) [42] and tensor decomposition (TD) [17, 43]. Here, we introduce a formulation with CP decomposition to solve the generalized space $(\bm{x})$ time $(t)$ parameter $(\bm{\theta})$ PDE, which is prohibitive for most numerical methods and machine learning approaches [42]. Consider a parameterized space-time PDE:
$\mathcal{L}u(\bm{x},t,\bm{\theta})=0,$ (10)
where $\mathcal{L}$ is the general partial differential operator (can be linear or nonlinear).
The INN solver uses the same neural network structure as the trainer, but the definition of loss function differs since training data is not known before solving the equation. Instead, the loss function is formulated as the weighted summation of the PDE residual [44, 45, 46]. INN approximates the general space $(\bm{x})$ time $(t)$ parameter $(\bm{\theta})$ solution using the CP decomposition format:
$\begin{split}\mathcal{J}{\bm{u}}(\bm{x},t,\bm{\theta};\bm{U},\bm{X})&=\sum_{m=1}^{M}[\mathcal{J}\prescript{m}{x_{1}}{\bm{u}}(x_{1})]\odot\cdots\odot[\mathcal{J}\prescript{m}{x_{d}}{\bm{u}}(x_{d})]\odot[\mathcal{J}\prescript{m}{t}{\bm{u}}(x_{t})]\\ &\odot[\mathcal{J}\prescript{m}{\theta_{1}}{\bm{u}}(\theta_{1})]\odot\cdots\odot[\mathcal{J}\prescript{m}{\theta_{k}}{\bm{u}}(\theta_{k})].\end{split}$ (11)
where $d$ and $k$ are the spatial dimension and the number of parameters, respectively. The operator $\odot$ refers to the element-wise multiplication. Similar to the trainer, the goal is to find nodal values $\bm{U}$ (and nodal coordinates $\bm{X}$, if one wants to adapt the mesh). As a result, INN obtains the space-time-parameter solution by minimizing the loss function:
@@
where $\delta\bm{u}$ is the weight function and can be defined using Galerkin, Petrov-Galerkin, collocation or quadratic formulation. It should be noted that INN should automatically satisfy Dirichlet and initial conditions. As a result, only weighted sum residual residual is considered in the loss function Eq. 11. See the Supplementary Information for detailed derivations of the loss function.
In Supplementary Information, a 3D linear elasticity problem is solved with the full interpolation INN and CP decomposition INN, and compared with FEM. Especially the INN with CP decomposition enables a voxel mesh of 10 million cube - $10^{21}$ grid points, which expands the state-of-the-art FEM limit by order of 10.
@@
where $\mathcal{J}\bm{u}(\bm{x})$ is the trained or solved forward model, $\bm{u}^{*}_{k}$ is the $k$-th measured observation and $\bm{x}^{*}$ is the calibrated input variable.
A few calibration benchmarks are provided in Supplementary Information that demonstrate the superior calibration accuracy and reliability of the INN calibrator.
3 Scientific and Engineering applications
@@
Figure 2: A summary of engineering applications of the INN methods to scientific and engineering problems. The calibration problem develops a reduced-order model of laser powder bed additive manufacturing to calibrate the heat source parameters from experimental data. A space-time-parameter solution of the same problem with large parameter space is demonstrated in Solving category. INN solvers result in significant storage reduction and fast simulation. Finally, a data-driven real-time online monitoring and feedback control tool is formulated with INN that uses only 18% training parameters and is 18-31 times faster compared to MLP.
@@
This section will illustrate how Engineering Software 2.0 enabled with INN can be applied to model various facets of metal additive manufacturing by integrating training, solving, and calibration processes. Note that the INN with CP decomposition will be referred to INN for brevity. Laser powder bed fusion (L-PBF) is a type of metal additive manufacturing in which metal powders are layered and melted by a laser beam to create objects that match the provided CAD designs [47]. The added flexibility of L-PBF comes with a very large design space and spatially varying microstructure after manufacturing. Therefore, considerable research is being devoted to employing in situ monitoring data for real-time control of L-PBF processes to minimize the variability [8]. Computational modeling of L-PBF spans the simulation of the manufacturing process with a high dimensional design space (solving), computing the internal relationship among different manufacturing parameters (calibrating), and designing an online monitoring and control system for L-PBF (training).
@@
The governing equation for modeling the manufacturing process is given by
$\frac{\partial T(\mathbf{x},t)}{\partial t}-k\Delta T(\mathbf{x},t)=Q_{h}(\mathbf{x},t,P,R,\eta,d)$ (14)
@@
This is a heat conduction equation where the moving laser heat source is incorporated through the source term, $Q_{h}$. The equation involves spatial parameter $\mathbf{x}$, temporal variable $t$, thermal conductivity $k$, power $P$, beam radius $R$, absorptivity $\eta$, and laser depth $d$. For the simplest case of single track laser pass, the initial condition is $T(\mathbf{x},0)=T_{0}(\mathbf{x})$, and the boundary conditions (BC) are $T=\tilde{T}$ in $\partial\Omega_{T}$ for Dirichlet BC and $\frac{\partial T}{\partial\mathbf{x}}\cdot\mathbf{n}=\bar{q}$ in $\partial\Omega_{q}$ for Neumann BC, where the spatial domain boundary $\partial\Omega=\partial\Omega_{T}\cup\partial\Omega_{q}$.
@@ 3.1 Calibration of Heat Source Parameters for Additive Manufacturing
In this example, the INN is used for calibrating a model of heat source of L-PBF, which uses a laser heat source to selectively melt metal powder bed to form a connected shape. A commonly used Gaussian beam profile can be modeled as a volumetric heat source [48]. The heat source equation reads,
$Q_{h}(\textbf{x},t,\textbf{p})=\frac{2\mathcal{P}\eta}{\pi r^{2}_{b}d}e^{-\frac{2(x^{2}+y^{2})}{r^{2}_{b}}}$ (15)
where, $z_{top}-z\leq d$, $z_{top}$ is z coordinate of the current powder bed surface. The notations follow Eq. 14 and $r_{b}$ is the standard deviation of the Gaussian source. To get an accurate representation of the physics behind it, model parameters need to be calibrated against experimental data. Following [49], three calibration parameters $p_{1}$, $p_{2}$, and $p_{3}$ can be defined as, $d=p_{1}\frac{\mathcal{P}}{V}RHF^{2}$, $\eta=p_{2}\frac{\mathcal{P}}{V}RHF^{2}$, and $r_{b}=p_{3}\frac{\mathcal{P}}{V}RHF^{2}$. Here, $P$ is laser power, $V$ is the scan speed, and $RHF$ is a residual heat measure predetermined based on a laser toolpath. The challenges for calibrating these parameters are: 1) the computational models are often very expensive to perform even one forward run, 2) the computational models are not automatically differentiable, and 3) limited experimental data is available. As a demonstration for this problem, INN with 1760 trainable parameters is trained on 44 sets of meltpool temperature data generated by the finite difference method (FDM) using 4 different power-scanning speed setup. Later, this trained INN is used for optimizing the three aforementioned parameters using gradient-based optimizers. To show the superiority of INN over traditional neural networks in such data-scarce applications, a 3-layer MLP with 2689 trainable parameters is trained using the same data fed to INN. Details regarding this calibration process are provided in the Supplementary Information.
Both INN-based and MLP-based calibrations are repeated 50 times to give a statistical comparison (see Figure 2). Even with only 44 sets of simulation data, the INN calibrated parameters that produce mean meltpool temperature (or TEP, thermal emission plank) within 7.9% of the experimental data. On the contrary, MLP produced a difference of 44.1% against experimental data, which shows that it hardly learns the underlying physics using such small data with even more trainable parameters.
@@ 3.2 Solving High Dimensional Space-Time-Parameter Heat Equation
The governing equation Eq.14 has 8 input variables; 3 for spatial domain ($x,y,z$), 1 for time domain ($t$), and 4 for parametric domain ($k,p,e,d$). See Supplementary Information for a detailed problem definition. This equation is not feasible for process design with most existing computational methods, including the data-driven method inspired by the finite element method (FEM) and the physics-informed neural network (PINN) [20].
To solve the problem with FEM-based data-driven method, one first needs to sample from the 4-dimensional parametric domain, followed by running the time-stepping-based finite element analysis for each sampled parameter set. Then the FEM solutions should be stored as offline training data, and a neural network surrogate model is trained using the data. The computational cost for such operations is prohibitive since large space-time mesh is required in each simulation for real engineering application, as detailed in the Supplementary Information. Moreover, as a larger number of training data is generated due to the high dimensionality of the input domain, this method becomes impractical in terms of data storage. Furthermore, running a large-scale FEM simulation requires a large amount of computational memory.
PINN is a unified framework of training and solving for such parameterized PDEs; however, the computational cost still grows exponentially as the number of input variables grows. As illustrated in Supplementary Information, it is very challenging to solve this 8-dimensional equation. In addition, the convergence of solution accuracy is not guaranteed especially for such high-dimensional problems. Thus, it remains to be seen if PINN can even find a good solution for more than 7 input variables. PINN also suffers from memory issues for GPU training due to the exponential growth of the collocation points although batch training can be adopted to alleviate memory usage.
Contrary to FEM and PINN, INN solves this space-time-parameter problem all-at-once. As it is based on the projected 1D interpolation, the number of trainable parameters increases linearly with the input variables, making the computational cost manageable. Meanwhile, since tensor-decomposed data structures are leveraged, INN has the benefits of significant memory efficiency and storage gain, especially for high-dimensional problems. In Figure 2, we solve the 8-dimensional space-time-parameter PDE with the total number of space-time-parameter grid points of $3\times 10^{32}$, where we use 25,000 grid points to discretize each spatial and parametric coordinate and 50 grid points for the temporal coordinate. The details of the solver example are provided in the Supplementary Information. Since the number of trainable parameters of INN grows linearly, INN has a significant storage gain of $1.7\times 10^{25}$ compared to the classical FEM.
Moreover, when using INN, once the parametric solution of the AM problem is obtained, one can easily investigate the uncertainty propagation by imposing probabilistic distributions on the inputs of INN, thanks to the tensor-decomposed data structure. Details on the statistical prediction of parametric solution can be found in the Supplementary Information.
In this application, the INN trainer is used to devise an online control system for L-PBF additive manufacturing. The goal is to maintain a homogeneous melt pool temperature across each layer. The hypothesis is that a homogeneous melt pool temperature across each layer of material deposition will reduce the variability in the microstructure and mechanical properties. The fundamental challenge of implementing a model predictive control system is the computational time required for the forward prediction (involving finite element or computational fluid dynamics models). For prediction, the data is generated by fusing experiments and computational methods (see Supplementary Information for details). The online control loop uses the trained INN model to inform the manufacturing machine. There are two models: forward model and inverse model.
The forward model uses the mechanistic features (precomputed based on the tool path and geometry of the build to generalize a data-driven model for unseen tool path and geometric features; more details can be found in [50]), and melt pool temperature of the previous layer (measured by the sensor as Thermal Emission Plank (TEP) signal; more details in the Supplementary Information) to predict the TEP/melt pool temperature in the next layer for an uncontrolled process. The uncontrolled TEP is converted to a homogenized TEP over the entire layer. The inverse model takes this targeted, homogenized TEP as the input along with the previous layer’s mechanistic features to predict the required laser power for the next layer. The laser power is then exported to the machine to print the next layer. The forward model is trained on experimentally observed data while the inverse model is trained on computational data coming from a finite element solver. The INN model is applied for both forward and inverse prediction and compared against MLP. Figure 2 shows that for the inverse model INN is at least 18 to 31 times faster to train compared to the MLP to reach the same level of training error, and it uses only about 18% of the parameters that MLP uses. The speed-up depends on the amount of data used. The INN method is consistently more accurate compared to MLP (see Supplementary Information for more details).

### v4

@@ 2 INN: Unification of Training/Learning, Solving, and Calibrating
A general INN forward propagation is provided in Eq.3 as a tensor contraction. The INN trainer optimizes nodal values $\bm{U}=\{\bm{u}^{(j)}\}_{j=1,\cdots,J}$ (and if needed, nodal coordinates ($\bm{X}=\{\bm{x}^{(j)}\}_{j=1,\cdots,J}$)) under a given loss function and training data. Consider a regression problem with $K$ labeled data: $(\bm{x}^{*}_{k},\bm{u}^{*}_{k})$, $k=1,\cdots,K$; and $\bm{x}^{*}_{k}\in\mathbb{R}^{I},\bm{u}^{*}_{k}\in\mathbb{R}^{L}$. The superscript * denotes data. A mean squared error (MSE) loss function for this regression problem is defined as:
@@
Finally, an optimization is formulated as:
@@
Section 3.1 exhibits the application of INN trainer to metal additive manufacturing while other benchmarks are elucidated in SI Section 2.
@@
An INN solver generalizes classical numerical methods such as FEM and meshfree, as well as the model order reduction methods such as proper generalized decomposition (PGD) [41] and tensor decomposition (TD) [14, 42]. Here, we introduce a formulation with CP decomposition to solve the generalized space $(\bm{x})$ time $(t)$ parameter $(\bm{\theta})$ (S-T-P) PDE, whose computational cost is prohibitively high for most numerical methods and machine learning approaches [41].
Consider a parameterized space-time PDE:
$\mathcal{L}\bm{u}(\bm{x},t,\bm{\theta})=0,$ (10)
where $\mathcal{L}$ is the general partial differential operator (can be linear or nonlinear). The INN solver uses the same neural network structure as the trainer, but the loss function varies with the equations to be solved. It is formulated as the weighted summation of the PDE residual [43, 44, 45]. An INN solution field in the S-T-P domain can be written as:
$\begin{split}\mathcal{J}{\bm{u}}(\bm{x},t,\bm{\theta};\bm{U},\bm{X})&=\sum_{m=1}^{M}[\mathcal{J}\prescript{m}{x_{1}}{\bm{u}}(x_{1})]\odot\cdots\odot[\mathcal{J}\prescript{m}{x_{d}}{\bm{u}}(x_{d})]\odot[\mathcal{J}\prescript{m}{t}{\bm{u}}(x_{t})]\\ &\odot[\mathcal{J}\prescript{m}{\theta_{1}}{\bm{u}}(\theta_{1})]\odot\cdots\odot[\mathcal{J}\prescript{m}{\theta_{k}}{\bm{u}}(\theta_{k})],\end{split}$ (11)
where $d$ and $k$ are the spatial dimension and the number of parameters, respectively. Similar to the trainer, the goal is to find nodal values $\bm{U}$ (and nodal coordinates $\bm{X}$, if one wants to adapt the mesh). As a result, the INN obtains the S-T-P solution by minimizing the loss function:
@@
where $\delta\bm{u}$ is the weight function and can be defined using Galerkin, Petrov-Galerkin, collocation or quadratic formulation. Due to the Kronecker delta property of INN interpolation functions, Dirichlet boundary conditions and initial conditions can be strongly imposed. As a result, only the weighted sum residual is considered in Eq. 11 as a loss function. This distinguishes INNs from most data-driven PDE solvers that weakly impose those conditions [46]. See the SI Section 4 for detailed derivations of the loss function and solution scheme. INN solvers applied to the metal additive manufacturing and the linear elastic solid mechanics problem are introduced in Section 3.3 and SI Section 4, respectively.
@@
where $\mathcal{J}\bm{u}(\bm{x})$ is the trained or solved forward model, $\bm{u}^{*}_{k}$ is the $k$-th measured observation and $\bm{x}^{*}$ is the calibrated input variable. The INN calibrator applied to the heat source calibration task in metal additive manufacturing can be found in Section 3.3 while other benchmarks can be found in SI Section 3.
3 Scientific and Engineering applications: Metal Additive Manufacturing
@@
Figure 3: A summary of engineering applications of INNs to additive manufacturing (AM) problems. First, a data-driven real-time online monitoring and feedback control tool is formulated with an INN that uses only 18% training parameters of MLP and is 18-31 times faster than MLP. The calibration problem develops a reduced-order model of laser powder bed fusion (L-PBF) AM to calibrate the heat source parameters from experimental data. Finally, the INN solver solves a space-time-parameter heat transfer equation, resulting in a significant storage reduction and faster simulations.
@@
This section illustrates how Engineering Software 2.0 enabled with INNs can be applied to advance various facets of metal additive manufacturing (AM). Note that the INN with CP decomposition will be referred to as INN in this section for brevity. Laser powder bed fusion (L-PBF) is a type of metal additive manufacturing in which metal powders are layered and melted by a laser beam to create objects that match the provided CAD designs [47]. The added flexibility of L-PBF comes with a huge design space and spatially varying microstructure after manufacturing. Therefore, considerable research is devoted to employing in situ monitoring data for real-time control of L-PBF processes to minimize the variability and optimize the target properties [48]. Computational modeling of L-PBF spans the simulation of the manufacturing process with a high dimensional design space (solving), the identification of model parameters from sparse experimental data (calibrating), and the online monitoring and control of the L-PBF (training).
@@
The governing equation for modeling the manufacturing process is given by a heat conduction equation with a moving laser heat source $S_{h}$:
$\frac{\partial T(\mathbf{x},t)}{\partial t}-k\Delta T(\mathbf{x},t)=S_{h}(\mathbf{x},t,P,\eta,d)$ (14)
@@
where the equation involves spatial variables $\mathbf{x}$, temporal variable $t$, and four variable parameters: thermal conductivity $k$, laser power $P$, absorptivity $\eta$, and laser depth $d$. For the simplest case of single track laser pass, the initial condition is $T(\mathbf{x},0)=T_{0}(\mathbf{x})$, and the boundary conditions (BCs) are $T=\tilde{T}$ in $\Gamma^{D}$ for Dirichlet BC and $\frac{\partial T}{\partial\mathbf{x}}\cdot\mathbf{n}=\bar{q}$ in $\Gamma^{N}$ for Neumann BC, where the spatial domain boundary $\Gamma=\Gamma^{D}\cup\Gamma^{N}$.
The first application is fully data-driven, where the data are generated from numerical methods and the INN trainer is used to develop an online control system for L-PBF AM. The goal is to maintain a homogeneous temperature of the melt pool (i.e., a molten region near the laser spot) across each layer using real-time sensor data, under the hypothesis that a homogeneous melt pool temperature across each layer of material deposition will reduce the variability in the microstructure and mechanical properties. The laser power at each point of a certain layer needs to be controlled (increased/decreased by certain amount compared to previous layer) to achieve this homogeneity in melt pool temperature. The fundamental challenge of implementing a model predictive control system for such an application is the computational resources required for forward prediction (involving finite element or computational fluid dynamics models) and inverse prediction, as illustrated in Figure 3. The INN trainer aims to provide a reliable and memory-efficient surrogate model that can replace mechanistic models in predictive control systems. In this application, the database is generated by fusing experiments and computational methods (see SI Section 2.6 for details) for aluminum alloy. The online control loop uses the trained INN models to inform the manufacturing machine. There are two models: forward model and inverse model.
The forward model uses the mechanistic features and the thermal emission plank (TEP, a representative of the melt pool temperature) to predict the TEP of the next layer of the build when no correction is applied. The mechanistic features are precomputed based on the tool path and geometry of the build to generalize a data-driven model for unseen tool path and geometric features [49]). The target TEP for the next layer is the average of the forward model’s outputs (i.e., predicted TEPs for the next layer) measured at the current layer. The inverse model takes this target TEP as the input along with the previous layer’s mechanistic features to predict the required laser power for the next layer. Finally, the control software calculates how much change in laser power is required at each point, and sends the information to the machine. The corrected laser power is then used to print the next layer.
The forward model is trained on experimentally observed data, while the inverse model is trained on computational data coming from a finite-difference solver [18]. The INN model is applied for both forward and inverse prediction and is compared with the MLP. Figure 3 shows that for the inverse model INN is at least 18 to 31 times faster to train compared to the MLP to reach the same level of training error, and it reduced the number of trainable parameters by a staggering amount of 82%. The speed-up depends on the amount of data used. The INN method is consistently more accurate compared to MLP (see, SI Section 2.6 for more details).
@@ 3.2 Calibration of Heat Source Parameters for Additive Manufacturing
In this example, an INN calibrates the heat source function $S_{h}$ in Eq.14 for experimental data. A Gaussian beam profile is modeled as a volumetric heat source [50], which is written as:
$S_{h}(\mathbf{x},t,\mathbf{p})=\frac{2P\eta}{\pi r^{2}_{b}d}e^{-\frac{2(x^{2}+y^{2})}{r^{2}_{b}}}\cdot\mathbf{1}_{\{z|z_{top}-z\leq d\}}(z)$ (15)
where $z_{top}$ is the z-coordinate of the current surface of the powder bed and $\mathbf{1}$ is the indicator function. The notation follows Eq. 14, and $r_{b}$ is the Gaussian profile’s standard deviation that controls the beam’s width. Following reference [10], the heat source parameters $d,\eta,r_{b}$ are controlled by the calibration parameters $p_{1}$, $p_{2}$, $p_{3}$ such that $d=p_{1}\frac{P}{V}RHF^{2}$, $\eta=p_{2}\frac{P}{V}RHF^{2}$, and $r_{b}=p_{3}\frac{P}{V}RHF^{2}$. Here, $V$ is the laser scan speed and $RHF$ is a predetermined residual heat factor based on the laser toolpath [51].
This calibration problem’s major challenge is achieving a highly accurate surrogate model given a sparse set of training data. INNs outperform MLPs in solving these data-scarce problems because they are rooted in the interpolation theory. As a demonstration, an INN with 1760 trainable parameters is trained on 28 sets of meltpool temperature data generated by the finite difference method (FDM) [18] using 4 different power-scanning speed set-ups. Later, this trained INN is used as a surrogate to optimize the three parameters ($p_{1},p_{2},p_{3}$) given the experimental data. To show the superiority of INNs over traditional neural networks in such data-scarce applications, a 3-layer MLP with 2689 trainable parameters (roughly the same order of magnitude as the INN) is trained using the same data fed to the INN. In practice, we observed that INN converges well within 100 epoch, while MLP requires over 1000 epoch to converge to a test MSE 10 times larger than INN given all other setup the same. Details on this training and calibration process are provided in SI Section 3.1.
The INN- and MLP-based calibrations are repeated 50 times to provide a statistical comparison (see Figure 3). Even with only 28 sets of simulation data, the calibrated parameters using the INN produce a mean melt pool temperature within a 7.9% difference of the experimental data. In contrast, the parameters calibrated from the MLP surrogate produce a difference of 44. 1%, which is unsatisfactory. This is expected since MLP is a general model requiring a large amount of data to achieve a reliable representation of the problem. On the other hand, the INN achieves a much lower difference, showing that the model can learn underlying physics even with a scarce dataset.
@@ 3.3 Solving High Dimensional Space-Time-Parameter Heat Equation
[…]
The surrogate model $f$ can also be obtained using the classical data-driven approach. In the offline stage, training data can be generated by repeatedly running numerical solvers (such as finite difference method (FDM) and variational multiscale method (VMS)) in the Space-Time domain with different sets of parameters $\mathbf{{p}}$. Then a surrogate model is trained with the data. The standard data-driven approach suffers from the curse of dimensionality when the parametric space is high-dimensional. As a result, the data-driven approaches face three challenges at the offline stage: expensive computation for repetitive data generation, huge memory cost for running full-scale simulations, and excessive disk storage of offline training data. We estimate the time and storage required for the data-driven approach with FDM and VMS, as shown in Table 2 and illustrated in Figure 2.
Unlike the data-driven approaches, INN solvers treat the parameters $\mathbf{p}$ as additional parametric inputs. As a result, INNs can directly obtain a parametric surrogate model $f=u(\mathbf{x},t,\mathbf{p})$ directly from the governing equation without going through the cumbersome offline data generation and surrogate model training process. This is particularly beneficial to high-dimensional parametric surrogate models. Moreover, our results prove the INN solution (or the S-T-P surrogate model) achieves an exceptionally high accuracy of $R^{2}$ = 0.9969, considering that most data-driven approaches for S-T-P problems trained on numerical simulation data suffer from low model accuracy, typically below $R^{2}=0.9$ [52]. It is worth mentioning that the physics-informed neural networks (PINNs) can handle the S-T-P problem similar to the INN solver. However, the number of collocation points scales exponentially with the number of inputs, and our preliminary study revealed that PINN fails to converge for the same 8-dimensional S-T-P problem after consuming considerable computational resources (see SI Section 4.4). The performance comparison of INN versus data-driven methods is shown in Table. 2.
Table 2: Performance comparison of different surrogate models. The INN’s discretization in the 4D parametric space is $100^{4}$, so the corresponding sampling points required for the data-driven methods should be $10^{8}$. The total simulation time and storage requirement of data-driven approaches are estimated from a single run. Detailed explanations for this benchmark can be found in SI Sections 4.3 and 4.4.
Method
Resolution
($\mu$m)
Number of
offline simulations
Total offline
simulation time
Storage
Data-driven (FDM) 25 $1\times 10^{8}$ 2,600 years $7.82\times 10^{8}$
Data-driven (VMS) 12.5 $1\times 10^{8}$ 9 years $8.79\times 10^{6}$
INN 8 $1$ 15 minutes $7.37\times 10^{-2}$